# Marharyta Domnich, Julius Välja, Rasmus Moorits Veski, Giacomo Magnifico, Kadi Tulver, Eduard Barbu, Raul Vicente

### v1

@@ Introduction
Evaluating counterfactual explanations poses a significant challenge in the field. While various quantitative metrics, such as validity, proximity, sparsity, coherence, robustness, and diversity (Guidotti 2022; Karimi et al. 2022; Rasouli and Chieh Yu 2024) are currently used, they often fall short in capturing the human perspective, missing key explanatory virtues and leading to inconsistent findings that complicate the development of a standardized evaluation framework. It is commonly recommended that user studies should be conducted to assess the efficacy of counterfactual explanations as ”excellent computational explanations may not be good psychological explanations” (Keane et al. 2021). Despite this, such studies are rarely utilized for benchmarking counterfactual explanations (Longo et al. 2024). One of the reasons for this is the difficulty and expense of recruiting a sufficient number of experts capable of performing these evaluations. Even when executed, user studies do not guarantee consistent and reproducible results as perceptions of what constitutes a reasonable explanation can vary widely between individuals and user groups (Kenny et al. 2021). Furthermore, most studies only employ a few qualitative measures, such as satisfaction and trust, which fail to address the nuanced features influencing human preferences (Warren, Byrne, and Keane 2023). While human assessments of counterfactual explanations are invaluable, these issues of cost and scalability make it very challenging to make meaningful comparisons and generalizations between multiple frameworks or domains.
@@
Recognizing the limitations of existing methodologies, this paper explores the potential of Large Language Models (LLMs) to serve as a benchmark for automating the evaluation of counterfactual explanations. Current LLMs have demonstrated remarkable capabilities in interacting with natural language data, from extensive data summarisation (Liu et al. 2024) and pattern deduction (Jin et al. 2024) to idea generation (Girotra et al. 2023) and problem-solving through branching solutions (Yang et al. 2024), and many more (Wang et al. 2024a). Based on these premises, LLMs are hypothesized to mimic human evaluative judgments effectively, offering a more accessible and cost-efficient alternative to traditional methods.
In light of these considerations, this paper addresses the following question: Can the evaluation process of counterfactual explanations be effectively automated using LLMs? To answer this question we created a diverse set of 30 counterfactual scenarios that were varied across different dimensions of explanatory qualities. The scenarios were evaluated by 206 human respondents in overall satisfaction and metrics of feasibility, consistency, completeness, trust, fairness, complexity and understandability. Next, we divided data for fine-tuning several LLM models to assess every metric score and compared the results to human data on a reserved test set. The pipeline can be seen in Figure  1.
@@
In the following section, we review user studies that focus on evaluating counterfactual explanations and the potential of LLMs in simulating human responses.
User studies in evaluating counterfactual explanations
In addition to quantitative explanatory metrics like proximity, validity, or sparsity, most researchers agree that it is crucial to also capture the subjective preferences of human users in aiming for more human-centric AI explanations (Kirsch 2017; Keane et al. 2021; Longo et al. 2024). Yet, a survey found only 21% of 100 studies on counterfactual methods included user evaluations (Keane et al. 2021). Furthermore, many of those studies test the use of counterfactual explanations vs no-explanations rather than comparing different methods, leaving only 7% of papers that report user evaluations for benchmarking different counterfactual algorithms.
In recent years, some user studies have been conducted with tabular counterfactual data. For instance, (Warren, Keane, and Byrne 2022) conducted a study with 127 participants to compare the effects of counterfactual and causal explanations on both objective prediction accuracy and subjective judgments of satisfaction and trust. (Bove et al. 2023) explored the impact of plural counterfactual examples on objective understanding and a modified version of the Explanation Satisfaction Scale (Hoffman et al. 2018) in a lab study with 112 participants. (Förster et al. 2021) conducted a study with 46 participants assessing the realism and typicality of an explanation. Two user studies have benchmarked counterfactual methods for perceived practicality of users in a study with 135 participants (Ghazimatin et al. 2020), and an online study with 500 responders (Spreitzer, Haned, and van der Linden 2022). Additionally, (Akula et al. 2022) tested their approach on image data, evaluating justified trust as quantitative metric and explanation satisfaction as qualitative metric across different algorithms.
Overall, user studies on explanation satisfaction often focus on a limited range of aspects (Mueller et al. 2019), typically measuring satisfaction and trust, while neglecting other essential qualities of the explanations themselves. These studies may not adequately capture human preferences, which are influenced by context, presentation, and cognitive biases, especially when preferences are not clearly defined (Covell 2019; Kliegr, Bahník, and Fürnkranz 2021; Tversky and Simonson 1993). As a result, studies that fail to capture the full spectrum of explanatory qualities contribute to a narrow and inconsistent perspective of human judgements, leaving a significant gap in our understanding of which features are central to good explanations.
@@
Predicting human evaluation using Machine Learning has garnered widespread acceptance in various domains, such as human-computer interaction (Kiseleva et al. 2016; Yang, Levow, and Meng 2012), recommendation systems (Siro, Aliannejadi, and De Rijke 2023), speech quality assessment (Reddy, Gopal, and Cutler 2022), etc. The progressive advancement of LLMs’ causal reasoning abilities (Bhattacharjee et al. 2024) suggests their usage in the context of explainability, since the explanations in natural language generated with these processes present qualities akin to those of human output (Castelnovo et al. 2024) and the explanatory process can be further enhanced through a post-output chat pipeline (Slack et al. 2023). LLMs have also been used to evaluate and model user satisfaction to provide insight regarding choices and preferences (Kim et al. 2024), to directly simulate user feedback for model tuning (Ebrat and Rueda 2024), and as artificial user / model-in-tuning pairs (Gao et al. 2024). However, to the best of our knowledge there is currently no work related to simulating human assessment in evaluating of counterfactual explanations with LLMs.
@@
Training LLMs to evaluate the quality of counterfactual explanations as humans do requires human-labeled data. As of the writing of this article, there exists no widely-used dataset of human-evaluated counterfactual explanations. To fill this gap, we created a varied dataset of 30 counterfactual explanation instances, which were graded on 8 different criteria by 206 people through an online survey.
@@
For selecting the dimensions to include in our study, we reviewed literature on qualitative metrics influential to human judgements. Among the most frequently cited explanatory virtues are coherence and simplicity (Mackonis 2013), aligning with the understanding of human mental models and a preference for consistent and parsimonious information (Johnson-Laird 2010). Coherence as a qualitative metric can be measured internally, representing consistency within the explanation, or externally, taking into account the prior knowledge of the rater (Zemla et al. 2017). Our work focuses on internal coherence to measure consistency between different parts of the explanation, independent of an individual’s prior experiences.
@@
The virtue of simplicity is also discussed under the terms (Desired) Complexity (Zemla et al. 2017) and Selection (Vilone and Longo 2021), assuming people prefer simple explanations (Lombrozo 2007). However, evidence suggests humans sometimes favor complex explanations involving more causal links (Zemla et al. 2017), or that moderate complexity and sufficient detail are preferred (Ramon et al. 2021; Hoffman et al. 2018). For this study, we chose to include the metric of Complexity, with desired values falling in the middle, as explanations can be perceived as either too simple or too complex.
@@
A commonly assessed quality in user studies is Trust. Various definitions focus on trust in the method generating explanations (Perrig, Scharowski, and Brühlmann 2023; Scharowski et al. 2024). Trust in explanations is considered in terms of trustworthiness, evaluating the perceived credibility of suggested changes (Stepin et al. 2022). We define Trust as belief that following the explanation would lead to the desired outcome.
@@
Feasibility is one of the most agreed-upon metrics when discussing counterfactual explanations, although discussed under different names: Controllability (Byrne 2019), Actionability (Rasouli and Chieh Yu 2024) and even split into Actionability and Mutability (Karimi et al. 2022). While actionability has also been employed as a quantitive measure (Guidotti 2022), feasibility refers to whether the proposed changes are perceived as achievable and realistic. Research indicates that explanations failing this criterion are rated poorly (McCloy and Byrne 2000; Butz et al. 2024).
@@
Understandability, also known as Readability (Stepin et al. 2022) or Comprehensibility (Ali et al. 2023; Vilone and Longo 2021), relates to how effectively an explanation conveys the model’s decision process to the user or how easily the user grasps it. Generally, higher understandability is linked to greater user satisfaction, with clear and comprehensible explanations generally preferred, though complex answers may be favored in some contexts.
@@
Completeness has previously been discussed as Incompleteness (Zemla et al. 2017) or Informativeness, the latter of which also includes the notion of extraneous information (Stepin et al. 2022), is tied to understanding causal relations and partially depends on domain knowledge (Keil 2006). Evaluating completeness is challenging as people tend to fill logical gaps in explanations (Strickland and Keil 2011).
@@
Finally, the dimension of Fairness in counterfactual explanations has also been highlighted in recent work (Wang et al. 2024b). Due to the concern of models unintentionally encoding or even amplifying biases present in training data (Corbett-Davies et al. 2023), it is crucial to address potential unfairness and discrimination. Fairness has mostly been viewed as a quantitative metric (Ge et al. 2022) with little understanding of how it influences the perceived quality of explanation.
@@
Relying on previous work on human preferences and explanatory virtues, we selected 8 different criteria capturing a range of relevant dimensions (see previous section for an overview) guiding the creation of diverse counterfactual scenarios. The Adult dataset (Becker and Kohavi 1996) and the Pima Indians Diabetes dataset (Bennett, Burch, and Miller 1971) were chosen as a basis when formulating the counterfactual explanations, as they encompass a varying level of domain knowledge and include both categorical and continuous data. To ensure the dataset consists of diverse counterfactual explanations, we included explanations constructed fully from the features of the datasets as well as explanations that were constructed from artificial data in the final set. All the counterfactual scenarios were designed from the perspective of improving the factual situation, as directionality has also been shown to influence how explanations are perceived (Kuhl, Artelt, and Hammer 2023).
@@
We aimed to include examples of explanations that fulfilled the different qualities at varying levels, in order to train LLM models to be able to distinguish between good and bad explanations. We included specific instances where different metrics had been varied, with the exception of Understandability and Overall satisfaction. We did not specifically vary the overall satisfaction of explanations, as this metric was to serve as a general indicator of the perceived quality of an explanation and as a benchmark for other metrics. Also, all explanations were designed to be as understandable as possible, and we did not include instances with purposefully poor wording to ensure that all respondents understood what they were reading, and could therefore reliably assess the other metrics.
@@
Our dataset contained examples of extreme changes in both categorical and continuous types of features, as it has been suggested that people may evaluate these differently (Warren, Byrne, and Keane 2023). For example, we explored how humans perceive Feasibility by creating explanations which changed inactionable features (e.g. age); features by different margins (a 1000€ pay increase vs 10 000€); continuous features outside and within distribution and starting from the value 0; ordinal features in the infeasible direction (e.g. lowering education level). For Consistency, we changed features that are widely considered to be connected (e.g. hours studied and average grade) in both covarying and conflicting directions, with both categorical and continuous features. Differences in Completeness were implemented by having sufficiently detailed explanations as well as explanations with obvious gaps. Furthermore, useful context was added to certain questions to ensure minimal necessary domain knowledge on the topic, the lack of which could influence perceptions of completeness. Variety in Trust was induced by having logical, solution-oriented explanations and those unlikely to bring about the desired change. To include cases of poor Fairness, some examples contained recommendations to change features widely considered controversial (e.g. gender, age). For varying Complexity, we included instances that might be perceived as too complex as well as too simple by having explanations with a different length and number of recommendations to similar problems. Here, we hypothesised that a desired level of Complexity lies in the middle, which is also reflected in the slightly different scale of measurement compared to other metrics. All selected metrics, along with their definitions and scales as presented in the questionnaire, are detailed in Table 1.
@@
To assess the overall suitability and comprehensibility of the compiled scenarios and evaluation metrics, a pilot study was conducted with 15 volunteers recruited among university students and colleagues. Feedback gathered during the pilot led to revisions in the wording of some metric descriptions. Additionally, the Coherency metric was renamed to Consistency and Bias was changed to Fairness to aid comprehension for the participants.
The final version containing 30 counterfactual scenarios was then shared on the Prolific platform and evaluated by a total of 206 respondents on the basis of the metrics in Table 1. All metrics were rated on a 6-point ordinal scale from 1 (lowest) to 6 (highest) with the exception of Complexity, which was rated on a 5-point scale from -2 (too simple) to 2 (too complex), where the rating 0 corresponded to desired complexity. The scenarios were presented to the participants in randomised order while the evaluation metrics were kept in the same order. All respondents had to be at least 18 years of age and fluent in English to participate. The full questionnaire is available in Supplementary Materials [^2] The public link will be added for camera-ready version. and one question example can be seen in Appendix A, Table A.1.
To detect fraudulent participants, a hidden attention check was included in the questionnaire. Responses were also analysed based on response time, average understandability score, a clustering of the respondents, and the uniformity of response patterns. Additionally, individual answers to 3 indicator questions were analysed. For example, if a participant rated an explanation recommending a change in place of birth as feasible, that respondent was flagged. Respondents failing in 3 aforementioned criteria were removed from further analyses, in total 10 respondents were removed.
@@
The correlation diagram in Figure 2 shows that all explanatory qualities significantly correlate with each other (p-value $<10^{-4}$, $\alpha$ = $\frac{0.05}{28}$), except between Complexity and Fairness. An analysis of questions involving varied fairness revealed they lacked overly complex explanations. The intercorrelated responses are therefore likely to reflect that humans grade the explanations as a whole, rating different metrics in the context of the entire scenario and other explanatory virtues. Notably, all metrics correlate positively with satisfaction, highlighting their importance for evaluating the overall quality of counterfactual explanations. Furthermore, reducing the 7 metrics’ scores (excluding Overall Satisfaction) to a 2-dimensional space using t-SNE, and coloring by Satisfaction, shows a distinct distribution correlating with overall satisfaction, detailed in Appendix A, Figure A.1.
@@
With the questionnaire data as the input dataset, we aimed to test and fine-tune Large Language Models for automated evaluation of counterfactual explanations. The models selected for this were Llama 3.1 Instruct, Llama 3 Instruct (Dubey et al. 2024) and GPT-4 (OpenAI 2023). GPT-4 was accessed via OpenAI API and the Llama models were fine-tuned on HPC clusters with NVIDIA Tesla A100 GPUs using the transformers library by Huggingface (Wolf et al. 2020). QLoRA, which relies on rank decomposition matrices and quantization, was used for reducing memory requirements during fine-tuning  (Dettmers et al. 2023).
@@
After gathering and filtering questionnaire responses, further data processing was needed to create a useful dataset. For each question-metric pair, we used the average response from 196 participants as the final value. Complexity, originally rated on a -2 to 2 scale, was linearly scaled to align with the 1 to 6 scale used for other metrics. To minimize scale effects and enhance generalizability, we consolidated all metric values into three distinct categories. Data analysis suggested that the differences between scores of 1 and 2, 3 and 4, and 5 and 6 could be effectively compressed. Subsequent analyses confirmed that three classes adequately predicted outcomes. Thus, we classified values below 3 as ”low,” values between 3 and 4 as ”medium,” and values above 4 as ”high.” These categories were deliberately balanced to ensure an equal distribution across the classes. With 30 questions and 8 metrics per question, this resulted in 240 instances of metric evaluation in total.
@@
Importantly, the instruction part of the prompt was taken from the questionnaire directly to ensure that the task reflects the gathered data, and all changes were made in what is known as a “system prompt”. For this task, the following system prompts were developed:
@@
The instruction or “user prompt” was adapted from the questionnaire, meaning it contained a factual-counterfactual pair from the questionnaire, alongside a modified metric evaluation question, such as ”Please rate as ’low’ (very unfeasible), ’medium’ or ’high’ (completely feasible), how feasible is this explanation:”. Consequently, each counterfactual explanation resulted in 8 instances, one for every metric under evaluation. The specific phrasing of all three system prompts can be found in Appendix B.
All of the prompts were tested using preliminary data from 100 participants and 4 LLMs, including Mistral-7B Instruct, Llama 2 7B Chat, and 8B and 70B versions of Llama 3 Instruct. Based on the results, which are available in the Appendix B, Table B.1, the baseline prompt was selected for all further experiments.
@@
The optimal fine-tuning hyperparameters for every model were discerned through extensive testing (see Appendix C, Table C.1). All models were fine-tuned using a completion-only data collator from Hugginface’s trl library (Werra et al. 2020) to improve the predictive performance of the models. With a typical language modelling data collator, the model would have learned to predict the question text as well, but this was unnecessary for the task at hand. Due to its proprietary nature, GPT-4 was not used for fine-tuning.
@@
Results in Table 3 show that LLMs possess some ability to evaluate counterfactual explanations even with zero-shot learning, with the GPT-4 model reaching 63% accuracy on metric split and Llama 3 70B Instruct reaching 59% accuracy on question split. All of the models tested surpassed the average accuracy one can achieve through random guessing, which is 33% in a three class prediction task. Fine-tuning improved accuracy scores significantly, with the Llama 3 70B Instruct model reaching an accuracy of 85% on the metric split and the recent but significantly smaller Llama 3.1 8B Instruct matching the result. For question split, the highest-performing model was Llama 3 70B Instruct, which after fine-tuning achieved 81% accuracy for three class prediction across 8 metrics.
@@
Confusion matrices revealed that after fine-tuning, the best-performing models for both splits (Figure 3b, Figure 3d) made no errors where ’low’ was classified as ’high’ and vice versa for any metric, which suggests a high-level understanding of the metrics. Table 4 illustrates the improvements in accuracy of each individual metric after fine-tuning, with notable gains especially in Completeness (improving from 33% to 83% and 75% for the metric and question split, respectively), Complexity (from 42% to 75% and 83%), and Understandability, which achieved perfect accuracy. Importantly, Satisfaction showed substantial improvements reaching 96% for metric split and 88% for question split. Feasibility and Trust remain challenging for prediction, largely because assessing the feasibility and outcomes of categorical changes is complex and often unclear as to whether it would bring the desired outcome.

### v2

# Towards Unifying Evaluation of Counterfactual Explanations: Leveraging Large Language Models for Human-Centric Assessments

@@ Introduction
Evaluating counterfactual explanations poses a significant challenge in the field. While various quantitative metrics, such as validity, proximity, sparsity, coherence, robustness, and diversity (Guidotti 2022; Karimi et al. 2022; Rasouli and Chieh Yu 2024) are currently used, they often fall short in capturing the human perspective, missing key explanatory virtues, and leading to inconsistent findings that complicate the development of a standardized evaluation framework. User studies are commonly recommended to assess the efficacy of counterfactual explanations, as ”excellent computational explanations may not be good psychological explanations” (Keane et al. 2021). Despite this, such studies are rarely utilized for benchmarking counterfactual explanations (Longo et al. 2024). One of the reasons for this is the difficulty and expense of recruiting a sufficient number of experts capable of performing these evaluations. Even when executed, user studies do not guarantee consistent and reproducible results as perceptions of what constitutes a reasonable explanation can vary widely between individuals and user groups (Kenny et al. 2021). Furthermore, most studies only employ a few qualitative measures, such as satisfaction and trust, which fail to address the nuanced features influencing human preferences (Warren, Byrne, and Keane 2023). While human assessments of counterfactual explanations are invaluable, these issues of cost and scalability make it very challenging to make meaningful comparisons and generalizations between multiple frameworks or domains.
@@
Recognizing the limitations of existing methodologies, this paper explores the potential of Large Language Models (LLMs) to serve as a benchmark for automating the evaluation of counterfactual explanations. Current LLMs have demonstrated remarkable capabilities in interacting with natural language data, ranging from extensive data summarisation (Liu et al. 2024) and pattern deduction (Jin et al. 2024) to idea generation (Girotra et al. 2023) and problem-solving through branching solutions (Yang et al. 2024). Based on these premises, LLMs are hypothesized to mimic human evaluative judgments effectively, offering a more accessible and cost-efficient alternative to traditional methods.
In light of these considerations, this paper addresses the following question: Can the evaluation process of counterfactual explanations be effectively automated using LLMs? To answer this question, we created a diverse set of 30 counterfactual scenarios that were designed to vary across multiple dimensions of explanatory qualities. The scenarios were evaluated by 206 human respondents in overall satisfaction and metrics of feasibility, consistency, completeness, trust, fairness, complexity and understandability. Next, we divided data for fine-tuning several LLM models to assess every metric score and compared the results to human data on a reserved test set. The pipeline can be seen in Figure 1.
@@
In the following section, we review user studies that focus on evaluating counterfactual explanations, and the potential of LLMs to simulate human responses.
User studies for evaluating counterfactual explanations
In addition to quantitative explanatory metrics like proximity, validity, or sparsity, most researchers agree that it is crucial to capture the subjective preferences of human users to achieve more human-centric AI explanations (Kirsch 2017; Keane et al. 2021; Longo et al. 2024). Yet, a survey found that only 21% of 100 studies on counterfactual methods included user evaluations (Keane et al. 2021). Furthermore, many of those studies test the use of counterfactual explanations vs no-explanations rather than comparing different methods, leaving only 7% of papers that report user evaluations for benchmarking different counterfactual algorithms.
In recent years, some user studies have been conducted with tabular counterfactual data. For instance, (Warren, Byrne, and Keane 2023) conducted a study with 127 participants to compare the effects of counterfactual and causal explanations on objective prediction accuracy and subjective judgments of satisfaction and trust. (Bove et al. 2023) explored the impact of plural counterfactual examples on objective understanding and a modified Explanation Satisfaction Scale (Hoffman et al. 2018) in a lab study with 112 participants. (Förster et al. 2021) conducted a study with 46 participants assessing explanation realism and typicality. Two user studies have benchmarked counterfactual methods for perceived practicality of users in a study with 135 participants (Ghazimatin et al. 2020), and an online study with 500 responders (Spreitzer, Haned, and van der Linden 2022). Additionally, (Akula et al. 2022) tested their approach on image data, evaluating justified trust as quantitative metric and explanation satisfaction as qualitative metric.
Overall, user studies on explanation satisfaction often focus on a limited range of aspects (Mueller et al. 2019), typically measuring satisfaction and trust, while neglecting other essential qualities of the explanations themselves. Such studies may fail to capture human preferences, which are shaped by context, presentation, and cognitive biases, especially when preferences are ill-defined (Kliegr, Bahník, and Fürnkranz 2021; Tversky and Simonson 1993). As a result, this limited scope leads to inconsistent perspectives on explanatory qualities, leaving a significant gap in understanding which features define good explanations.
@@
Predicting human evaluation with machine learning has gained widespread acceptance in various domains, such as human-computer interaction (Kiseleva et al. 2016), recommendation systems (Siro, Aliannejadi, and De Rijke 2023), speech quality assessment (Reddy, Gopal, and Cutler 2022), etc. The progressive advancement of LLMs’ causal reasoning abilities (Bhattacharjee et al. 2024) suggests their usage in the context of explainability, as their natural language explanations exhibit qualities similar to human output (Castelnovo et al. 2024) and the explanatory process can be further enhanced through a post-output chat pipeline (Slack et al. 2023). LLMs have also been used to evaluate and model user satisfaction, providing insights into choices and preferences (Kim et al. 2024), and as artificial user-model tuning pairs (Gao et al. 2024).
At the time of this paper’s submission, no prior work existed on simulating human assessments for evaluating counterfactual explanations using LLMs. However, a concurrent effort has since emerged, replicating an existing user study that compared counterfactual and causal explanations by replacing human participants with seven LLMs (debona2024evaluatingexplanationsllmstraditional), demonstrating that LLMs can replicate some conclusions from the original study. Nonetheless, to the best of our knowledge, no work has yet focused on evaluating the intrinsic quality of counterfactual explanations.
@@
Training LLMs to evaluate the quality of counterfactual explanations in a human-like manner requires human-labeled data. Currently, no widely-used dataset of human-evaluated counterfactual explanations exists. To fill this gap, we created a varied dataset of 30 counterfactual explanation instances, which were graded on 8 different criteria by 206 people through an online survey.
@@
To select the dimensions for our study, we reviewed literature on qualitative metrics that influence human judgments. Among the most frequently cited explanatory virtues are coherence and simplicity (Mackonis 2013), aligning with the understanding of human mental models and a preference for consistent and parsimonious information (Johnson-Laird 2010).
Coherence can be measured internally, representing consistency within the explanation, or externally, taking into account the prior knowledge of the rater (Zemla et al. 2017). Our work focuses on internal coherence, measuring consistency within different parts of the explanation, independent of an individual’s prior experiences.
@@
The virtue of simplicity, also referred to as (Desired) Complexity (Zemla et al. 2017) or Selection (Vilone and Longo 2021), assumes that people prefer simple explanations (Lombrozo 2007). However, evidence suggests humans sometimes favor complex explanations involving more causal links, or that moderate complexity and sufficient detail are preferred (Zemla et al. 2017; Hoffman et al. 2018). In this study, we include Complexity as a metric, with desired values lying in the middle, as explanations may be perceived as overly simple or complex.
@@
A commonly assessed quality in user studies is Trust. Definitions often focus on trust in the system generating explanations (Perrig, Scharowski, and Brühlmann 2023). Trust in explanations is considered in terms of trustworthiness, or the perceived credibility of suggested changes (Stepin et al. 2022). We define Trust as the belief that following the explanation will lead to the desired outcome.
@@
Feasibility is one of the most agreed-upon metrics when discussing counterfactual explanations, although discussed under different names: Controllability (Byrne 2019), Actionability (Rasouli and Chieh Yu 2024) and split into Actionability and Mutability (Karimi et al. 2022). While actionability has been employed as a quantitative measure (Guidotti 2022), feasibility refers to whether the proposed changes are perceived as achievable and realistic. Explanations failing this criterion are rated poorly (Butz et al. 2024).
@@
Understandability, also known as Readability (Stepin et al. 2022) or Comprehensibility (Vilone and Longo 2021), relates to how effectively an explanation conveys the model’s decision process to the user and how easily it is grasped. Higher understandability is generally linked to greater user satisfaction, with clear explanations preferred, though complex answers may be favored in some contexts.
@@
Completeness previously discussed as Incompleteness (Zemla et al. 2017) or Informativeness, the latter of which also includes the notion of extraneous information (Stepin et al. 2022), is tied to understanding causal relations and partially depends on domain knowledge (Keil 2006). Evaluating completeness is challenging since people often fill logical gaps in explanations (Strickland and Keil 2011).
@@
Finally, the dimension of Fairness in counterfactual explanations has been highlighted in recent work (Wang et al. 2024b). Concerns about models unintentionally encoding or amplifying biases in training data (Corbett-Davies et al. 2023) make it crucial to address potential unfairness and discrimination. Fairness is mostly been viewed as a quantitative metric (Ge et al. 2022), with limited understanding of its influence on perceived explanation quality.
@@
Relying on previous work on human preferences and explanatory virtues, we selected 8 different criteria capturing a range of relevant dimensions (see previous section for an overview) to guide the creation of diverse counterfactual scenarios. The scenarios in our study are grounded in actual outputs from counterfactual algorithms applied to commonly used datasets for counterfactual explanation evaluation, such as the Adult dataset and the Pima Indians Diabetes dataset. In some cases, we modified the algorithmic outputs to ensure a broader representation across explanatory quality metrics. Additionally, we tailored certain features to enhance clarity for human evaluators based on feedback from a pilot study. All counterfactual scenarios were designed from the perspective of improving the factual situation, as directionality has been shown to influence how explanations are perceived (Kuhl, Artelt, and Hammer 2023).
@@
We included examples of explanations that fulfilled the different qualities at varying levels to train LLM models to distinguish between good and bad explanations. Specific instances were created by varying metrics, with the exception of Understandability and Overall satisfaction. We did not specifically vary the overall satisfaction of explanations, as this metric serves as a general indicator of the perceived quality of an explanation. Also, all explanations were designed to be as understandable as possible, and no instances with purposefully poor wording were included to ensure participants could reliably assess other metrics.
@@
Our dataset contained examples of extreme changes in both categorical and continuous features, as people may evaluate these differently (Warren, Byrne, and Keane 2023). For example, we explored how humans perceive Feasibility by creating explanations which changed inactionable features (e.g. age); features by different margins (a 1000€ pay increase vs 10 000€); continuous features outside and within distribution, starting from the value 0; and ordinal features in infeasible directions (e.g. lowering education level). For Consistency, we changed features widely considered connected (e.g. hours studied and average grade) in both covarying and conflicting directions, using categorical and continuous features. Differences in Completeness were implemented with sufficiently detailed explanations or those containing obvious gaps. Furthermore, useful context was provided for some questions to ensure minimal domain knowledge was sufficient, the lack of which could influence perceptions of completeness. Variety in Trust was induced by having logical, solution-oriented explanations and others unlikely to bring about the desired change. Poor Fairness was represented by recommendations involving controversial features (e.g. gender, age). To vary Complexity, we included instances that might be perceived as too complex as well as too simple by having explanations with a different length and number of recommendations to similar problems. Here, we hypothesised that a desired level of Complexity lies in the middle, which is also reflected in the slightly different scale of measurement compared to other metrics. All selected metrics, along with their definitions and scales as presented in the questionnaire, are detailed in Table 1.
@@
To assess the suitability and comprehensibility of the compiled scenarios and evaluation metrics, a pilot study was conducted with 15 volunteers recruited from university students and colleagues. Feedback gathered during the pilot led to revisions in the wording of some metric descriptions. Additionally, the Coherency metric was renamed to Consistency and Bias was changed to Fairness to aid comprehension for the participants.
The final version containing 30 counterfactual scenarios was shared on the Prolific platform and evaluated by 206 respondents. On average, completing the questionnaire took 42 minutes. All metrics were rated on scales detailed in Table 1, with a 6-point ordinal scale from 1 (lowest) to 6 (highest) except for Complexity, rated on a 5-point scale from -2 (too simple) to 2 (too complex), where 0 corresponded to desired complexity. Counterfactual scenarios were presented to the participants in randomised order, while the evaluation metrics remained in the same order. All respondents had to be at least 18 years of age and fluent in English to participate. The evaluated quationnaire is available on HuggingFace [^1] https://huggingface.co/datasets/anitera/CounterEval, and an example question is shown in Appendix A, Table A.1.
To detect fraudulent participants, a hidden attention check question was included in the questionnaire. Responses were also analyzed based on response time, average understandability score, respondence clustering, and the uniformity of response patterns. Additionally, individual answers to 3 indicator questions were reviewed. For example, if a participant rated an explanation recommending a change in place of birth as feasible, that respondent was flagged. Respondents failing the three aforementioned criteria were excluded from further analysis, with a total of 10 respondents removed.
@@
The correlation diagram in Figure 2 shows that all explanatory qualities significantly correlate with each other (p-value $<10^{-4}$, $\alpha$ = $\frac{0.05}{28}$), except for Complexity and Fairness. An analysis of questions involving varied fairness revealed they did not include overly complex explanations. The intercorrelated responses are likely to reflect that humans grade the explanations as a whole, rating different metrics in the context of the entire scenario and other explanatory virtues. Notably, all metrics correlate positively with satisfaction, highlighting their importance for evaluating the overall quality of counterfactual explanations. Furthermore, reducing the 7 metrics’ scores (excluding Overall Satisfaction) to a two-dimensional space using t-SNE, and coloring by Satisfaction, shows a distinct distribution correlating with overall satisfaction, as detailed in Appendix A, Figure A.1.
@@
With the questionnaire data as the input dataset, we aimed to test and fine-tune LLMs for automated evaluation of counterfactual explanations. The models selected for this were Llama 3.1 Instruct, Llama 3 Instruct (Dubey et al. 2024) and GPT-4 (OpenAI 2023). Llama models were fine-tuned on HPC clusters equipped with NVIDIA Tesla A100 GPUs using the transformers library by Huggingface (Wolf et al. 2020). QLoRA, which relies on rank decomposition matrices and quantization, was used to reduce memory requirements during fine-tuning  (Dettmers et al. 2023).
@@
After gathering and filtering questionnaire responses, further data processing was needed. The preprocessing scripts and model weights are publicly available in the GitHub repository [^2] https://github.com/anitera/CounterEval. For each question-metric pair, the average response from 196 participants was used as the final value. Complexity, originally rated on a -2 to 2 scale, was linearly scaled to align with the 1 to 6 scale used for other metrics. To minimize scale effects and enhance generalizability, we consolidated all metric values into three distinct categories: ”low,” ”medium,” and ”high.”. Data analysis suggested that the differences between scores of 1 and 2, 3 and 4, and 5 and 6 could be effectively compressed into these categories. This classification ensured a balanced distribution across the dataset. With 30 questions and 8 metrics per question, this resulted in 240 instances of metric evaluation in total.
@@
Importantly, the instruction part of the prompt was taken from the questionnaire directly to ensure that the task reflects the gathered data, and all changes were made in what is known as a “system prompt”. The following system prompts were developed:
@@
The instruction or “user prompt” was adapted from the questionnaire, meaning it contained a factual-counterfactual pair from the questionnaire, alongside a modified metric evaluation question, such as ”Please rate as ’low’ (very unfeasible), ’medium’ or ’high’ (completely feasible), how feasible is this explanation:”. Consequently, each counterfactual explanation resulted in 8 instances, one for every metric under evaluation. The specific phrasing of all prompts can be found in Appendix B.
All of the prompts were tested using preliminary data from 100 participants and four LLMs, including Mistral-7B Instruct, Llama 2 7B Chat, and 8B and 70B versions of Llama 3 Instruct. Based on the results ( Appendix B, Table B.1), the baseline prompt was selected for all further experiments.
@@
The optimal fine-tuning hyperparameters for every model were discerned through extensive testing (see Appendix C, Table C.1). All models were fine-tuned using a completion-only data collator from Huggingface’s trl library (Werra et al. 2020) to improve the predictive performance of the models. With a typical language modelling data collator, the model would have learned to predict the question text as well, but this was unnecessary for the task at hand. Due to its proprietary nature, GPT-4 was not used for fine-tuning.
@@
Results in Table 3 show that LLMs possess some ability to evaluate counterfactual explanations even with zero-shot learning, with the GPT-4 reaching 63% accuracy on metric split and Llama 3 70B Instruct reaching 59% on question split. All models surpassed the 33% accuracy expected from random guessing in a three-class task. Fine-tuning significantly improved scores, with the Llama 3 70B Instruct reaching 85% accuracy on the metric split and the recent but significantly smaller Llama 3.1 8B Instruct matching the result. For question split, the highest-performing model was Llama 3 70B Instruct, which after fine-tuning achieved 81% accuracy for three class prediction across 8 metrics.
@@
The confusion matrices show that after fine-tuning, the best-performing models for both splits (Figure 3b, Figure 3d) made no errors misclassifying ’low’ as ’high’ or vice versa, suggesting a high-level understanding of the metrics. Table 4 highlights accuracy improvements after fine-tuning, with notable gains in Completeness (improving from 33% to 83% and 75% for the metric and question split, respectively), Complexity (from 42% to 75% and 83%), and Understandability, which achieved perfect accuracy. Importantly, Satisfaction showed substantial improvements reaching 96% for metric split and 88% for question split. Feasibility and Trust remain challenging for prediction, largely because assessing the feasibility and outcomes of categorical changes is complex and often unclear as to whether it would bring the desired outcome.